# Visual Representation Learning Guided by Multi-modal Prior Knowledge

### v1

@@ Results
We conduct our few-shot learning tasks in the domain of road sign recognition. We first train the models with the GTSRB dataset. Then, we keep the encoder network and replace the decoder to fit the right class dimension in the target dataset. The new decoder is randomly initialized. The models are retrained with different amounts of labeled target data in the CTSD or the RTSD. We perform one-shot and five-shot learning experiments. One-shot learning involves using just one image per class from the target dataset for retraining. Similarly, five-shot learning uses five images per class in the target dataset for retraining. As Table 5 shows, our KGV model achieves 7.0% performance gain in the CTSD dataset and 5.2% in the RTSD dataset in one-shot learning. In a five-shot learning setting, the KGV model achieves 0.2 % improvement in CTSD and 2.3% improvement in RTSD. The overall results of our few-shot learning experiments indicate that our KGV model is able to adapt significantly better to the new data distribution.
@@
In this paper, we introduce KGV, a knowledge-guided visual representation learning method that leverages multi-modal prior knowledge to improve generalization in deep learning. The multi-modal prior knowledge across data distributions in KGV helps to mitigate overfitting the training dataset, thereby enhancing the robustness of the model against data distribution shifts. We observe that incorporating additional visual priors, such as shapes, colors, and sketches, guided by factual knowledge from the KG, is crucial for KGV’s performance. However, we also observe that using these visual priors on baseline models, such as ResNet50, without guidance from the KG may have little or even a negative impact. Further, modeling node embeddings of the KG as Gaussian embeddings enhances the performance of KGV compared to traditional translation-based KGE methods. Even in scenarios where the training and test distributions have no significant shift, our KGV method still demonstrates notable improvements. Additionally, our KGV method exhibited significant data efficiency in the low data regimes.
@@
A main objective of our future work is to scale our concept to larger datasets and KGs. Additionally, we plan to explore further methods for aligning KGEs and image embeddings, and inducing information of the KG at inference time. Lastly, we are also interested in whether incorporating additional text information will improve the performance.
Appendix
The following sections provide detailed information about the datasets and the training process for each scenario in our experiments. Additionally, we present further analysis through ablation experiments.
Dataset Settings
Detailed descriptions of the used datasets are provided in this section.
Road Sign:
Road SignInformativeMandatoryProhibitoryNo PedestrianSpeed Limit 30Warning Warning Pedestrian DE CN RS Warning Animal Road Sign Feature ShapeCircleTriangleSign LegendExclamationTruckColorRedWhite
Figure 4: An abstract of hierarchical relations existing in the road sign recognition domain.
We leverage three commonly used road sign recognition datasets to evaluate our model. They are the German Traffic Sign Recognition Benchmark (GTSRB) from Stallkamp et al. (2011), Chinese Traffic Sign Dataset (CTSD) from Yang et al. (2015), and Russian Traffic Sign Images Dataset (RTSD) from Shakhuro and Konouchine (2016). The detailed information on these three datasets and the synthetic images of road sign elements are listed in Table 6 and Table 7. The hierarchical relations are shown in Figure 4.
Dataset Total Images Road Sign Classes
GTSRB 51970 43
CTSD 6164 58
RTSD 32983 64
Table 6: GTSRB, CTSD, and RTSD represent three road sign recognition datasets of China, Germany, and Russia. The total images and class numbers are listed in the table.
Feature Types Sub-class Image Sub-class Image
Black [Uncaptioned image] Blue [Uncaptioned image]
Color Brown [Uncaptioned image] Green [Uncaptioned image]
Red [Uncaptioned image] White [Uncaptioned image]
Yellow [Uncaptioned image]
Circle [Uncaptioned image] Diamond [Uncaptioned image]
Shape Octagon [Uncaptioned image] Triangle-Down [Uncaptioned image]
Triangle-Up [Uncaptioned image]
Animal [Uncaptioned image] Bicycle [Uncaptioned image]
Exclamation [Uncaptioned image] Person [Uncaptioned image]
Sign Legend Sand [Uncaptioned image] SnowFlake [Uncaptioned image]
Truck [Uncaptioned image] TwoPerson [Uncaptioned image]
Vehicle [Uncaptioned image]
Table 7: Synthetic Images in the Road Sign Recognition Domain. Seven colors, five shapes, and nine sign legends are included. For each type, we generated synthetic images as illustrated in the table.
DVM-CAR:
Feature Types Sub-class Image Sub-class Image
Beige [Uncaptioned image] Black [Uncaptioned image]
Blue [Uncaptioned image] Bronze [Uncaptioned image]
Brown [Uncaptioned image] Burgundy [Uncaptioned image]
Gold [Uncaptioned image] Green [Uncaptioned image]
Grey [Uncaptioned image] Indigo [Uncaptioned image]
Color Magenta [Uncaptioned image] Maroon [Uncaptioned image]
Navy [Uncaptioned image] Orange [Uncaptioned image]
Pink [Uncaptioned image] Purple [Uncaptioned image]
Red [Uncaptioned image] Silver [Uncaptioned image]
Turquoise [Uncaptioned image] White [Uncaptioned image]
Yellow [Uncaptioned image]
Convertible [Uncaptioned image] Coupe [Uncaptioned image]
Estate [Uncaptioned image] Hatchback [Uncaptioned image]
Bodytype Limousine [Uncaptioned image] Minibus [Uncaptioned image]
MPV [Uncaptioned image] Pickup [Uncaptioned image]
Saloon [Uncaptioned image] SUV [Uncaptioned image]
Van [Uncaptioned image]
Table 8: Synthetic Images in the Car Recognition Domain. Twenty-one colors and eleven body type sketches are included. We generate synthetic images for each color and collect body type sketches on the internet.
Deep Visual Marketing Car (DVM-CAR) dataset (Huang et al. 2021) is created from 335,562 used car advertisements. It contains 1,451,784 images of cars from various angles, as well as their sales and technical data. We use this dataset to predict the car model from the given images. All images are resized to the resolution of 128 $\times$ 128. We use the same training, validation, and testing setting as MMCL (Hager, Menten, and Rueckert 2023). Their dataset consists of 286 classes with a total of 418,100 training samples, 104,308 validation samples, and 525,557 testing samples. Samples of synthetic images used for training can be seen in Table 8. We generate 20 pure color images with small variations for each color category. For each body type category, we collect 20 sketch images. Note that we extract the image instance-level triplets from the tabular annotations provided in the DVM-CAR dataset. Figure 5 shows an abstract of relations between one sample image instance and the other nodes in the KG.
Refer to captionColorSilverRedBody TypeCoupeSUVGear TypeAutoManuelFuel TypeDieselPetrolNumber25Viewpoint045315hasColorhasBodyTypehasDoorNohasViewpointhasFuelTypehasGearTypePorsche 911instanceOf
Figure 5: The relations between one image instance of Porsche 911 with other nodes in the car recognition KG.
ImageNet:
ThingFoodBeverageMilkBeerEatCornHotdogPlaceArchitectureBuildingBuildingBookshopDomeSpaceNaturalSpaceCliffCoralReefSpeciesArchaeaBacteriaEukaryoteAnimalBirdMammaCatDogBoxerSalukiFungusPlantToolHairSlideDeviceCameraInstrumentOrganVehicleAirplaneShipAircraftCarrierYawl
Figure 6: An abstract of the hierarchical relations existing in the ImageNet domain.
Mini-ImageNet (Vinyals et al. 2016) is a subset of the ImageNet dataset, consisting of 60K images of size $84\times 84$ with 100 classes, each having 600 examples. ImageNet-Skecth (Wang et al. 2019) comprises 50000 images, 50 images for each of the 1000 ImageNet classes. Here, we collect the 100 classes in mini-ImageNet from 1000 classes of ImageNet-Sketch for training. We employ three distinct datasets for our evaluation: ImageNetV2 (Recht et al. 2019), ImageNet-R (Hendrycks et al. 2021a), and ImageNet-A (Hendrycks et al. 2021b). ImageNetV2 comprises 10 fresh test images for each class, sharing the same set of labels as mini-ImageNet. ImageNet-R includes various artistic representations such as art, cartoons, and deviant styles, covering 200 classes from ImageNet. Lastly, ImageNet-A features unaltered, real-world examples that occur naturally. For our purposes, we assess only those classes that overlap with mini-ImageNet. An abstract of hierarchical relations extracted from the WordNet in the ImageNet Domain is illustrated in Figure 6.
Training Details
Our baseline ResNet50 has a simple structure of a ResNet50 backbone as an encoder and full connect layers as a decoder. The ResNet50 backbone is untrained and randomly initialized in a Xavier manner. Our KGV model and the other baselines, GCNZ and DGP, use the same image encoder backbone. The node and relation lookup tables, which contain the embeddings of nodes and relations in the knowledge graph, are also initialized, following the initialization method proposed in transE (Bordes et al. 2013). The relation, represented as ‘sub class of/instance of’ is an exception. It is initialized with a vector of all zeros and frozen during the training procedure. We use one NVIDIA A100 GPU with 80GB memory to train the models. Each algorithm is run with three random seeds to get the average performance.
The hyperparameters used in the three domains are illustrated as follows:
  • Road Sign Recognition Domain: The weight balance $\beta$ defined is 0.1. We train the baseline and our model with the Adam optimizer and a constant learning rate of 0.0002. The total training epochs are 120. The mini-batch size is 32. We resize all the input images to $32\times 32$.
  • Car Recognition Domain: We resize all input image to 128 $\times$ 128. The batch size we select is 512. Adam optimizer with a learning rate of 0.001 and cosines annealing learning rate schedule is utilized for optimization. The training epochs for 5%, 10%, 50%, and 100% of data are 600, 200, 150, and 100, respectively. The balance $\beta$ we use is 0.5.
  • ImageNet Domain: Both Baseline and our model KGV are trained with Adam optimizer and a learning rate of 0.001. The batch size we select is 1024 and the training epoch is 150. We resize all the input images to $84\times 84$. The balance $\beta$ we use is 0.5.
Additional Analysis
In this section, we further analyze which knowledge plays the most important role in the performance. We conduct ablation experiments on road sign recognition, car recognition and ImageNet domains.
Road Sign Recognition:
As Table 9 demonstrates, we conducted ablations studies for our KGV method, excluding one type of prior knowledge among sign legends, colors, and shapes. Our findings indicate that prior knowledge of shapes is the most critical factor in enhancing performance under the distribution shift from German to Chinese and Russian road signs. This is likely because road signs with the same semantic meaning share identical shapes in Germany, China, and Russia. Prior knowledge of sign legends also significantly contributes to performance, while information about colors is the least impactful.
Model CTSD RTSD
KGV 74.8% 79.3%
KGV w/o sign legend 73.4% (-1.4%) 75.5% (-3.8%)
KGV w/o colors 73.7% (-1.1%) 77.2% (-2.1%)
KGV w/o shapes 72.8% (-2.0%) 74.7% (-4.6%)
Table 9: Classification Accuracy in the Road Sign Recognition Domain. The models are trained with 43 classes data of GTSRB and testing with shared 25 classes data of CTSD and shared 36 classes data of RTSD.
Car Recognition:
We conducted ablation experiments on the DVM-CAR dataset. These experiments were performed using a 5% training data setting, as our KGV method exhibited the largest performance margin compared to the baselines in this scenario. As Table 10 shows, we found that prior knowledge of the body type is the most important factor in improving performance. Without this information, there is a significant drop of 4.0%. Additionally, prior knowledge of colors is also crucial, as the performance drops by 2.4% without color information.
Model Top-1 Accuracy
ResNet50 64.0%
KGV 77.1%
KGV w/o body type 73.1% (-4.0%)
KGV w/o color 74.7% (-2.4%)
KGV w/o viewpoint 76.2% (-0.9%)
KGV w/o non-morpho. 75.2% (-1.9%)
Table 10: Top-1 Accuracy in DVM-CAR Model Classification with 5% Training Data.
ImageNet:
Model KGV KGV w/o color KGV w/o sketch
Mini-ImageNet 66.8% 63.6% -3.2% 64.4% -2.4%
ImageNetV2 52.7% 50.0% -2.7% 51.5% -1.2%
ImageNet-R 41.4% 39.5% -1.9% 29.7% -11.7%
ImageNet-A 6.8% 5.1% -1.7% 6.2% -0.6%
Table 11: Classification Accuracy in the ImageNet Domain.
As demonstrated in Table 11, we performed ablation studies on our KGV method by excluding either prior knowledge of color or sketch. Our results indicate that knowledge of color significantly impacts performance on the Mini-ImageNet, ImageNetV2, and ImageNet-A datasets. The knowledge of sketch proves to be more critical for classifying the ImageNet-R dataset, likely due to the similarity between the ImageNet-Sketch and ImageNet-R datasets.

### v2

@@ Results
Prior Knowledge Model Road Sign Domain ImageNet Domain Car Domain
Imp. KG GTSRB CTSD RTSD Mini-ImageNet ImageNetV2 ImageNet-R ImageNet-A DVM-CAR
P-ResNet50+KGV 99.5% 74.5% 79.8% 81.8% 62.4% 56.6% 6.4% 97.0 %
P-ResNet50 99.3% (+0.2%) 73.1% (+1.4%) 72.5% (+7.3%) 75.0% (+6.8%) 61.3% (+1.1%) 51.6% (+5.0%) 4.4% (+2.0%) 96.2% (+0.8%)
DINOv2+KGV 79.1% 31.2% 62.4% 95.4% 92.4% 87.0% 85.4% 92.8%
DINOV2+LP 78.7% (+0.4%) 28.9% (+2.3%) 59.9% (+2.5%) 94.9% (+0.6%) 91.5% (+0.9%) 86.1% (+0.9%) 84.3% (+1.1%) 89.3% (+3.5%)
CLIP-I+KGV 89.3% 73.0% 67.8% 90.1% 89.1% 83.2% 50.1% 90.5%
CLIP-I+LP 89.1% (+0.2%) 73.5% (-0.5%) 60.0% (+7.8%) 89.7% (+0.4%) 88.9% (+0.2%) 82.9% (+0.3%) 49.6% (+0.5%) 89.2% (+1.3%)
Table 6: Pre-trained ResNet50, DINOv2 (1B), and CLIP Image encoder, which are denoted as P-ResNet50, DINOv2, and CLIP-I, are used for our experiments. LP represents linear probing for fine-tuning. Our KGV’s encoder is replaced by one of these pre-trained image encoders denoted by X+KGV.
Model Road Sign Domain ImageNet Domain Car Domain
GTSRB CTSD RTSD Mini-ImageNet ImageNetV2 ImageNet-R ImageNet-A DVM-CAR
KGV 99.3% 74.8% 80.1% 66.8% 52.7% 41.4% 6.8% 96.3%
KGV w/o Gaussian Emb. 99.4% (-0.1%) 74.7% (+0.1%) 79.9% (+0.2%) 61.2% (+5.6%) 49.0% (+3.7%) 37.8% (+3.6%) 4.8% (+2.0%) 96.3%(+0.0%)
KGV w/ TransH 99.3% (0.0%) 73.5% (+1.3%) 79.3% (+0.0%) 61.8% (+5.0%) 48.2% (+4.5%) 38.6% (+2.8%) 5.5% (+1.3%) 95.8% (+0.5%)
KGV w/ DistMult 99.2% (+0.1%) 73.7% (+1.1%) 75.6% (+5.4%) 60.5% (+6.3%) 47.6% (+5.1%) 37.5% (+3.1%) 4.7% (+2.1%) 95.1% (+1.2%)
KGV w/ RGCN 99.3% (+0.0%) 74.1% (+0.7%) 78.5% (+1.6%) 63.5% (+3.3%) 50.6% (+2.1%) 39.7% (+1.7%) 6.1% (+0.7%) -
Table 7: KGV with alternative score function designs like TransH, DistMult, and replacing Gaussian embeddings with vector embeddings are listed in the table. KGV with GNN-based method like RGCN are also compared.
@@
We conduct our few-shot learning tasks in the domain of road sign recognition. We first train the models with the GTSRB dataset. Then, we keep the encoder network and replace the decoder to fit the right class dimension in the target dataset. The new decoder is randomly initialized. The models are retrained with different amounts of labeled target data in the CTSD or the RTSD. We perform one-shot and five-shot learning experiments. One-shot learning involves using just one image per class from the target dataset for retraining. Similarly, five-shot learning uses five images per class in the target dataset for retraining. As Table 5 shows that KGV achieves 7.0% performance gain in the CTSD dataset and 5.2% in the RTSD dataset in one-shot learning. In a five-shot learning setting, the KGV model achieves 0.2 % improvement in CTSD and 2.3% improvement in RTSD.
Combining Implicit and Explicit Prior Knowledge
In this sub-section, we investigate if combining implicit prior knowledge from the pre-trained models and explicit multi-modal prior knowledge from KG and visual prior images can further improve our method’s performance. We conduct experiments by replacing the random initialized ResNet50 encoder with pre-trained ResNet50, DINOv2(1B), and the image encoder of CLIP. As shown in Table 6, applying our method to these pre-trained encoders consistently enhances performance across all three domains. Although DINOv2 and CLIP have already achieved strong results on the ImageNet domain, the incorporation of KGV further improves their performance. These results indicate that our method can be adapted to state-of-the-art pre-trained image encoders to further enhance their performance.
Ablation Study
Score Functions
Here, we compare four score function designs: (1) replacing the Gaussian-form node embedding with vector embedding (TransE), (2) using TransH’s score function, (3) adopting DistMult’s score function, and (4) incorporating R-GCN, a GNN-based method. As Table 7 shows, overall KGV performs best among all these score function designs. The performance of KGV without Gaussian embedding drops significantly in the ImageNet domain compared to other domains. This highlights the importance of using Gaussian embedding for node representation to handle inclusion relations between object categories and their respective images. The DistMult demonstrates suboptimal performance across all three visual domains. While GNN-based method R-GCN outperform TransH and DistMult, it still does not surpass the effectiveness of our original score function design.
Knowledge types
Here, we further analyze the impact of relations types in the KGV’s performance. We conduct ablation experiments on road sign recognition, car recognition.
As Table 8 demonstrates, in each study, one type of relation is excluded, i.e. sign legends, colors, and shapes. Our findings indicate that shape is the most critical factor in enhancing performance under the distribution shift from German to Chinese and Russian road signs. This is likely because road signs with the same semantic meaning share identical shapes in Germany, China, and Russia. Sign legends also significantly contributes to performance, while information about colors is the least impactful.
Model CTSD RTSD
KGV 74.8% 80.1%
KGV w/o sign legend 73.4% (+1.4%) 75.5% (+4.6%)
KGV w/o colors 73.7% (+1.1%) 77.2% (+2.9%)
KGV w/o shapes 72.8% (+2.0%) 74.7% (+5.4%)
Table 8: Classification Accuracy in the Road Sign Domain. The models are trained with 43 classes data of GTSRB and testing with shared 25 classes data of CTSD and shared 36 classes data of RTSD.
On the DVM-CAR dataset, we performed experiments in a setting with 5% training data. We observe that even with this amount of data, KGV exhibits a large performance margin compared to the baselines. As indicated in Table 9, the information about the body type is the most important factor in improving performance. Without this information, there is a significant drop of 4.0%. Additionally, prior knowledge of colors is also crucial, as the performance drops by 2.4% without color information.
Model Accuracy Model Accuracy
KGV 77.1% ResNet50 64.0% (+13.1%)
KGV w/o body type 73.1% (+4.0%) KGV w/o color 74.7% (+2.4%)
KGV w/o viewpoint 76.2% (+0.9%) KGV w/o non-morpho. 75.2% (+1.9%)
Table 9: Accuracy in DVM-CAR Model Classification with 5% Training Data.
@@
In this paper, we introduce KGV, a knowledge-guided visual representation learning method that leverages multi-modal prior knowledge to improve generalization in deep learning. The multi-modal prior knowledge across data distributions in KGV helps to mitigate overfitting the training dataset, thereby enhancing the robustness of the model against data distribution shifts. We observe that modeling node embeddings of the KG as Gaussian embeddings enhances the performance of KGV to model hierachical relations. Furthermore, our method demonstrates a strong tendency to be data efficient, i.e consistently achieves good performance in the low data regimes. Finally, we show that our KGV method can be well adapted to the SOTA visual foundation models such as DINOv2 and CLIP.
@@
A main objective of our future work is to scale our concept to larger datasets and KGs. Additionally, we plan to explore further methods for aligning KGEs and image embeddings, and inducing information of the KG at inference time. Lastly, we will look whether incorporating additional text information will improve the performance.